Barbara bought a pair of dress shoes and a pair of sports shoes. The combined cost was $90.48. The dress shoes cost $15.88 less than the sport shoes. Find the cost of each pair of shoes.
Answer by checkley77(12844)

x+x+15.88=90.48
2x=90.48-15.88
2x=74.60
x=74.60/2
x=$37.30 was the price of the sport shoes.
37.30+15.88=$53.16 was the price of the dress shoes.
Proof:
37.30+53.18=90.48
90.48=90.48